OpenStudy (kirbykirby):

A red box has 4 balls and 6 black balls. A black box has 6 red balls and 4 red balls. Step 1: Toss a coin. If heads: pick a ball from the red box. If tails: pick a ball from the black box. Step 2: Choose a ball from the box having the same colour as the ball selected in step 1. Let X = number of red balls you draw in the 1st two steps. Find E(X).

OpenStudy (kirbykirby):

The solution used indicator variables (sorry the dashes in the cases, i don't know how to add space in LaTeX).R = red, B= black \( I_1 = \begin{cases} 1 & ball-in-step1-is-red\\ 0 & otherwise \end{cases} \) \( I_2 = \begin{cases} 1 & ball-in-step2-is-red\\ 0 & otherwise \end{cases} \) \(X=I_1+I_2\) \(E(X)=E(I_1)+E(I_2)\) \(E(X)=P(I_1=1)+P(I_2=1)\) They get \(P(I_1=1)=0.5\) which I understood, but for \(P(I_2=1)\): \(P(I_2=1)=P(2^{nd} ball=R)\) \(P(I_2=1)=P(2^{nd}=R|1^{st}=R)*P(1^{st}=R)+P(2^{nd}=R|1^{st}=B)*P(1^{st}=B)\) \(=(4/10)*(1/2)+(6/10)*(1/2)\) I'm having so much trouble how they get 4/10 for P(2nd=R|1st=R)... if you picked a red ball initially from the red box won't the probability be less than 4/10?

OpenStudy (kirbykirby):

The probalem doesn't state anything though if balls are being picked with or without replacement

OpenStudy (anonymous):

then you can't do it

OpenStudy (kirbykirby):

are they doing it correctly if there's assuming with replacement? which is what I think they're doing

OpenStudy (kirbykirby):

they're assuming*